A person places a cup of coffee on the roof of her car while she dashes back into the house for a forgotten item. When she returns to the car, she hops in and takes off with the coffee cup still on the roof.

. Part A If the coefficient of static friction between the coffee cup and the roof of the car is 0.29, what is the maximum acceleration the car can have without causing the cup to slide? Ignore the effects of air resistance. Express your answer using two significant figures.

Part B What is the smallest amount of time in which the person can accelerate the car from rest to 17 m/s and still keep the coffee cup on the roof? Express your answer using two significant figures.
Answer #1

A)
In the vertical:
Fnety = 0
N - mg = 0
N = mg

In the horizontal:
Fnet = ma
Friction = ma
Uk(N) = ma
Uk(mg) = ma
Ukg = a
0.29*9.8 = a
a =2.8 m/s^2

*the minimum acceleration to keep it up top

B)
a = (vf - vi) / time
2.842 = (17- 0) / time
t = 5.98 s